OpenStudy (anonymous):

Volumes of Revolution help. Find the volume formed by rotating the region enclosed by: y=x^2, x=y^2, about the line x=-2

OpenStudy (tkhunny):

Find your points of intersection so you know the limits of your integration.

OpenStudy (anonymous):

Would they be 0 and 1?

OpenStudy (tkhunny):

Perfect. Decide, Shells or Disks? It's about the same difficulty on this nice region. Just pick one.

OpenStudy (tkhunny):

Are you struggling with the decision? Really, just flip a coin. Maybe both ways?

OpenStudy (anonymous):

Shell than the outter equation would be x^2 +2 and inner sqrt(x) +2?

OpenStudy (anonymous):

i got the +2 just becuase of the x=-2 thats its flipping around, i dont know if its right

OpenStudy (tkhunny):

Shells it is, then... \(2\pi\int\limits_{0}^{1} Radius\cdot Height\; d(Radius)\) "outer" and "inner" don't mean the same thing for Shells. You just need to subtract them to get the height of each shell.

OpenStudy (tkhunny):

The "Radius" inside the integral is simply (-2-x). Do you see this? This gives: \(2\pi\int\limits_{0}^{1}(-2-x)(Height)\;dx\)

OpenStudy (anonymous):

yeah this is making a lot of sense actually

OpenStudy (tkhunny):

Excellent. All we've left is the height of each shell. What say you?

OpenStudy (anonymous):

so would that be the difference of sqrtx and x^2?

OpenStudy (tkhunny):

That's correct! Get them in the right order.

OpenStudy (anonymous):

so its (x^2 -sqrtx) and than the integral of that times the radius?

OpenStudy (tkhunny):

You have it! Let's see what you get.

OpenStudy (anonymous):

2pi(2-(1/2)(1/3)(1-2(1)^(3/2)))

OpenStudy (anonymous):

its not right though

OpenStudy (tkhunny):

Hmmm... What did your original integrand look like? It should be \((-2-x)(x^{2}-\sqrt{x})\). Did we get that far?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

That expands to a mess: \(2x^{1/2} - x^{3} - 2x^{2} + x^{3/2}\) Did you get all that?

OpenStudy (anonymous):

yeah let me try it again real quick i may have made a little mistake

OpenStudy (anonymous):

i integrated it to this (2 x^(5/2))/5+(4 x^(3/2))/3-x^3 than plugged in 1 and multiplied it by 2pi. Its still wrong though

OpenStudy (tkhunny):

What's up with "-x^3"? You should have -(1/4)x^4 - (2/3)x^3

OpenStudy (anonymous):

-2pi((2 (1)^(5/2))/5+(4 (1)^(3/2))/3-(1/4)(1)^4 - (2/3)(1)^3) Is this what its supposed to look like?

OpenStudy (anonymous):

without the - in the front

OpenStudy (anonymous):

Its right, i was off with my integration, thank you a ton for your help!

OpenStudy (tkhunny):

That looks like it!

OpenStudy (tkhunny):

It's a lot to wade through. GREAT WORK!!